Β Consider the following differential equation system which describes the interaction between two species.


π‘₯Μ‡= π‘₯(βˆ’0.7 + 1.6𝑦)

𝑦̇ = 1.3𝑦(1 βˆ’ 𝑦) βˆ’ 0.5π‘₯𝑦


a) What type of interaction is modeled by the system? What is the significance of the variables?

b) Find the equilibrium points and analyze their stability.Β 


1
Expert's answer
2021-05-23T16:43:20-0400

The interaction type is a competing exception. Variables make sense as values of mutually exclusive goods, or goods.


"\frac {\delta X}{\delta x}=-0.7+1.6y=0"

"\frac{\delta Y}{\delta y}=1.3-2.6y-0.5x=0"

"y=0.44"

"x=0.312"